Find the quadratic function that fits the data points for(-1,0) (3,0) (5,3)
Answer by stanbon(75887)

Find the quadratic function that fits the data points for(-1,0) (3,0) (5,3)
------------------------
Form::
ax^2 + bx + c = y
-----
Using (-1,0) you get a - b + c = 0
Using (3,0) you get 9a + 3b + c = 0
Using (5,3) you get 25a + 5a + c = 3
--------------------------------------------
Use any method you know to solve the system:
a = 1/4
b = -1/2
c = -3/4
-----
Equation:
y = (1/4)x^2 -(1/2)x -(3/4)
===================================
Cheers,
Stan H.